A piece of zinc was added to 1000cm3 of 0.2M hydrochloric acid. After effervescence had stopped, 28cm3 of the resulting solution required 17cm3 of 0.08M sodium trioxocarbonate IV solution for complete neutralization. Calculate the mass of the zinc added
1
Expert's answer
2020-06-19T07:00:59-0400

Sodium trioxocarbonate (IV) is a name for sodium carbonate, Na2CO3.

Ctotal(HCl) = 0.2 M

Vtotal(HCl) = 1000 cm3 = 1 L

C(Na2CO3) = 0.08 M

V(Na2CO3) = 17 cm3 = 0.017 L


Solution:

Balanced chemical reactions:

(1): Zn(s) + 2HCl(aq) = ZnCl2(aq) + H2(g)

(2): Na2CO3(aq) + 2HCl(aq) = 2NaCl(aq) + CO2(g) + H2O(l)


Determine the concentration of excess HCl using equation (2):

According to the equation (2): n(Na2CO3) = n(HCl)/2

Thus:

2 × C(Na2CO3) × V(Na2CO3) = Cexcess(HCl) × Vexcess(HCl)

Cexcess(HCl) = 2 × C(Na2CO3) × V(Na2CO3) / Vexcess(HCl)

Cexcess(HCl) = 2 × (0.08 M) × (0.017 L) / (0.028 L) = 0.097 M


Determine the amount of HCl that reacts with zinc:

n(HCl) = [Ctotal(HCl) - Cexcess(HCl)] × Vtotal(HCl)

n(HCl) = (0.2 M - 0.097 M) × 1L = 0.103 mol


Determine the amount of Zn using equation (1):

According to the equation (1): n(Zn) = n(HCl)/2

Thus:

n(Zn) = (0.103 mol) / 2 = 0.0515 mol


Calculate the mass of Zinc:

m(Zn) = n(Zn) × M(Zn) = (0.0515 mol) × (65 g/mol) = 3.3475 g = 3.35 g

The mass of Zn is 3.35 g


Answer: The mass of Zn is 3.35 g